# Recovery of rational functions via Hankel pencil method and sensitivities of the poles

### v2

@@ Abstract
In the paper, we develop a new method for the recovery of rational functions. Our idea is based on the property that Fourier coefficients of rational functions have the exponential structure and reconstruction of this exponential structure with the ESPRIT method in the frequency domain. Further we present sensitivity analysis for poles of rational functions reconstructed with our method in case of unstructured and structured perturbations. Finally, we consider several numerical experiments and, using sensitivities, explain the recovery errors for poles.
@@ 1 Introduction
For $M\in\mathbb{N}$, we define a rational function $r(z)$ of type $(M-1,M)$ by means of its partial fraction decomposition
@@
where $\gamma_{j},\,z_{j}\in\mathbb{C}$, $\gamma_{j}\neq 0$ and poles $z_{j}$ are pairwise distinct.
@@
We study the question of recovery of rational functions (1.1) using as input data a finite set of their Fourier coefficients $\widehat{r}_{k}$ defined by (4.1). We use the property that Fourier coefficients of rational functions have the exponential structure and we reconstruct this exponential structure in the frequency domain applying the ESPRIT method [19, 17]. The idea of recovery of rational functions by estimation of signal parameters in the frequency domain is not new. Similar technique for the recovery of or approximation by rational functions was employed in [7, 20, 21, 22]. The paper [21] is of special interest to us. In [21], the author considered recovery of the rational function (1.1) based on using noisy input data and discussed reconstruction results for the poles $z_{j}$, $j=1,...,M$. It was pointed out that for different levels of noise, not all poles $z_{j}$ are reconstructed with the same accuracy, and sometimes even the locations of certain poles are not reconstructed correctly. From these numerical observations, we can conclude that some poles $z_{j}$ of the rational function (1.1) are more ”sensitive“ to perturbations in the input data than the others. Our goal in this paper is to present theoretical explanations of these numerical observations by applying the sensitivity analysis. The method developed in [21] employees the AAK theory [16] for the reconstruction of the exponential structure of the Fourier coefficients $\widehat{r}_{k}$. We modify these ideas, by applying the ESPRIT method in the frequency domain. Since ESPRIT is based on using the Hankel pencil, it makes our method suitable for the sensitivity analysis.
@@
In [23], the authors developed the sensitivity analysis for the eigenvalues of Loewner pencil with respect to the unstructured and structured perturbations. We referee the interested reader to [1] for further application of Loewner framework for model reduction of dynamical systems. Following approach in [23], we present the sensitivity analysis for the poles of $r$ in (1.1) computed by our method. According to our idea, poles located inside and outside the unit circle are reconstructed separately as eigenvalues of some Hankel matrix pencils. Further, we consider two kinds of Hankel pencil perturbations. For the sensitivity $\rho$ with respect to unstructured perturbations, we present the connection with the condition numbers of Vandermonde matrices that are used in factorizations of the corresponding Hankel matrices. The sensitivity $\eta$ with respect to structured perturbations of Hankel pencil is studied in [24]. We present modification of these results for our special rational reconstruction problem (1.1). Note that the theory of sensitivities of eigenvalues of Hankel pencil perturbations was considered also in [11, 3]. In [11] and [3], the authors measured both structured and unstructured perturbations by the same quantity. Similarly as in [23], we introduce two different quantities for structured $\eta$ and unstructured $\rho$ sensitivities, respectively, of poles of the rational function (1.1). By checking if $\rho$ or $\eta$ is ”large“ for some pole $z_{j}$, we measure the ill-disposedness of $z_{j}$, while consideration of $\eta$ gives us also the additional information regarding that to perturbation of which measurement $\widehat{r}_{k}$ the pole $z_{j}$ is the most sensitive.
@@
The method developed in this paper is also closely related to our ESPIRA method for the exponential recovery [8][10]. ESPIRA uses the property that the classical Fourier coefficients of exponential sums have a rational structure and reconstructs this rational structure with the AAA method for rational approximation [14] in the frequency domain. Thus, the idea of rational reconstruction used in this paper can be considered as converse to the idea of ESPIRA. We want to pay a special attention also to the paper [5], where authors developed a method for exponential reconstruction using the sensitivity analysis of Hankel pencil eigenvalues.
@@
Finally, we want to mention the other methods for the recovery of rational functions such as the AAA method [14], the RKFIT (Rational Krylov Fitting) method [4], the Vector Fitting algorithm [13] and the Loewner framework method [1]. In [12], the authors discussed all these algorithms for the recovery of rational functions using as input information given noisy set of samples of $r$. Since we are interested in the sensitivity analysis of the poles, we have to assume that we apply perturbations to the Fourier coefficients $\widehat{r}_{k}$. Note that we can introduce the corresponding modification of the method based on employing function value of $r$ by computing $\widehat{r}_{k}$ via the quadrature formula for some grid of points (see Remark 4.1). Modifying the algorithm in this way will allow us to compare its performance with known rational methods. Numerical experiments show us that our approach gives equally good and for some examples even better recovery results from given noisy function values than the AAA method.
@@
Notation. Let us agree that throughout the paper, we will use the matrix notation $\mathbf{X}_{L,K}:=(X_{ij})_{i,j=1}^{L,K}$ for complex matrices of size $L\times K$ and the submatrix notation $\mathbf{X}_{L,K}(m:n,k:\ell)$, to denote a submatrix of $\mathbf{X}_{L,K}$ with rows indexed $m$ to $n$ and columns indexed $k$ to $\ell$, where first row and first column has index 1 (even though the row and column indices for the definition of the matrix may start with 0). For square matrices, we often use the short notation $\mathbf{X}_{K}$ instead of $\mathbf{X}_{K,K}$. By $\mathbf{X}_{L,K}^{\dagger}$ we denote the Moore-Penrose inverse of $\mathbf{X}_{L,K}$ and by $\mathbf{X}_{K}^{-1}$ inverse of the invertible matrix $\mathbf{X}_{K}$. We consider $\|\mathbf{X}_{L,K}\|_{2}:=\sigma_{\max}(\mathbf{X}_{L,K})$ the 2-norm or spectral norm, and $\|\mathbf{X}_{L,K}\|_{\mathrm{F}}:=\left(\sum\limits_{i=1}^{L}\sum\limits_{j=1}^{K}|X_{ij}|^{2}\right)^{1/2}=\left(\sum\limits_{i=1}^{\min\{L,K\}}\sigma_{i}^{2}\right)^{1/2}$ the Frobenius norm of the matrix $\mathbf{X}_{L,K}$. The $\ell_{1}$- and $\ell_{2}$-norm of the vector $\boldsymbol{x}=(x_{1},\ldots,x_{K})$ are given as usually by $\|\boldsymbol{x}\|_{1}=\sum\limits_{j=1}^{K}|x_{j}|$ and $\|\boldsymbol{x}\|_{2}=\left(\sum\limits_{j=1}^{K}|x_{j}|^{2}\right)^{1/2}$ respectively.
@@ 2 ESPRIT method for the recovery of exponential sums
As was described in Introduction, we will apply the ESPRIT method [19, 17] to reconstruct the exponential structure of the Fourier coefficients of rational functions (1.1). In this section, we explain the main ideas of this method.
Let $M\in\mathbb{N}$ and $f$ be an exponential sum defined by
$f(t):=\sum\limits_{j=1}^{M}\gamma_{j}z_{j}^{t},\ \gamma_{j},z_{j}\in\mathbb{C},\ \gamma_{j}\neq 0,$ (2.1)
where nodes $z_{j}$ are pairwise distinct. Let $L,N\in\mathbb{N}$ be given such that $M\leq L\leq N$, $L$ is the upper bound for $M$. We want to reconstruct an exponential sum (2.1), i.e. to reconstruct parameters $M$, $\gamma_{j}$ and $z_{j}$ for $j=1,\ldots,M$, using as input data samples $\boldsymbol{f}:=\big{(}f(0),f(1),\ldots,f(2N-1)\big{)}^{T}$.
@@
Since Vandermonde matrices in (2.2) have full column ranks, the matrix $\mathbf{H}_{2N-L,L+1}$ has rank $M$. Therefore in case of exact given data, the order $M$ of the exponential sum (2.1) is reconstructed as the numerical rank of the matrix $\mathbf{H}_{2N-L,L+1}$.
@@
From the last property, we conclude that nodes $z_{j}$, $j=1,\ldots,M$ in (2.1) are eigenvalues of the matrix pencil (2.3). Let us denote by $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ the vector of eigenvalues of (2.3). In case when $M$ is known, we can choose $N=L=M$ and compute $z_{j}$ as eigenvalues of
@@
Now we discuss how we solve the matrix pencil (2.3). Consider the SVD of ${\mathbf{H}}_{2N-L,L+1}$,
@@
where ${\mathbf{U}}_{2N-L}$ and ${\mathbf{W}}_{L+1}$ are unitary square matrices and, in the case of unperturbed data,
@@
where $\sigma_{1}\geq\ldots\geq\sigma_{M}$ denote the positive singular values of ${\mathbf{H}}_{2N-L,L+1}$. Thus (2.7) can be simplified to
@@
where the (last) $L+1-M$ zero-columns in ${\mathbf{D}}_{2N-L,L+1}$ and the last $L+1-M$ rows in ${\mathbf{W}}_{L+1}$ are removed. Similarly as before, the SVDs of ${\mathbf{H}}_{2N-L,L}(0)$ and ${\mathbf{H}}_{2N-L,L}(1)$ can be directly derived from (2.8) and the matrix pencil in (2.3) now reads
@@
Thus, eigenvalues of the matrix pencil (2.3) can be computed as eigenvalues of the matrix pencil (2.9). Then we define ${\mathbf{W}}_{M,L}(0)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,1\!:\!L)$, ${\mathbf{W}}_{M,L}(1)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,2\!:\!L+1)$, and determine the vector $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ as vector of eigenvalues of $\left({\mathbf{W}}_{M,L}(0)^{T}\right)^{\dagger}{\mathbf{W}}_{M,L}(1)^{T}$.
Having nodes $z_{j}$, we reconstruct coefficients $\gamma_{j}$, $j=1,\ldots,M$ as a least squares solution of the system of linear equations
@@ 3 Sensitivities of Hankel pencil eigenvalues
The concept of sensitivities arises from the perturbation of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [11] and further considered in [3]. In [23], the notion of sensitivities was defined for Loewner matrix pencil. In the paper [23], the authors define two sensitivities, which correspond to structured and unstructured perturbations. Using similar ideas, in [24], the author considered the structured sensitivity of Hankel matrix pencil. In [5], the sensitivity analysis of Hankel pencil was applied to develop a new method for exponential analysis for the recovery from noisy data. In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3) and further in Sunbsection 4.2 we present modification of the results from this section to develop sensitivity analysis for the poles of rational functions.
3.1 Sensitivity of Hankel pencil eigenvalues with unstructured perturbations
Following approach offered in [23] for Loewner pencil, we will show that in case of unstructured perturbations of Hankel pencil, the sensitivity of eigenvalues depends on the numerical conditions of Vandermonde matrices from factorizations (2.4) and (2.5) (for $N=L=M$). We consider the generalized eigenvalue problem (GEP) for the Hankel matrix pencil (2.6),
@@ 3.1 Sensitivity of Hankel pencil eigenvalues with unstructured perturbations
where $\boldsymbol{p}=(p_{1},\ldots,p_{M})^{T}$ and $\boldsymbol{q}=(q_{1},\ldots,q_{M})^{T}$ are right and left eigenvectors of the GEP (3.1) respectively. Then under the perturbation
$\tilde{\mathbf{H}}_{M}(0):=\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)},\ \ \tilde{\mathbf{H}}_{M}(1):=\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)},$
@@
Formula (3.2) is obtained from the perturbed system
@@
Assume that the norm of perturbation matrices is bounded as follows
@@
where $\varepsilon>0$. Then from (3.2) taking into account (3.3), we get for the GEP (3.1)
@@
where $\rho$ is the sensitivity of the eigenvalue $z$ of the GEP (3.1) defined by
@@
We derive further formulas for the exact computation and upper bounds of sensitivities $\rho_{j}$ of eigenvalues $z_{j}$, $j=1,\ldots,M$ of the GEP (3.1). Since Hankel matrices $\mathbf{H}_{M}(0)$ and $\mathbf{H}_{M}(1)$ satisfy factorizations (2.4) and (2.5) for $N=L=M$, left $\boldsymbol{p}^{(j)}$ and right $\boldsymbol{q}^{(j)}$ eigenvectors of the GEP (3.1) corresponding to the eigenvalue $z_{j}$ are given by formulas
@@
where $\boldsymbol{e}_{j}$ is the $j$th canonical vector in $\mathbb{R}^{M}$. Then using again the factorization (2.4) of the Hankel matrix $\mathbf{H}_{M}(0)$, we get
@@
From (3.4) taking into account (3.6), we obtain that the sensitivity $\rho_{j}$ of the eigenvalue $z_{j}$ is given by
@@
The ill-disposedness of the eigenvalue $z_{j}$ with respect to unstructured perturbations can be measured by checking whether $\rho_{j}$ is ”large“. Applying formulas (2.4), (2.5) and (3.5) to (3.7), we obtain for $j=1,\ldots,M$
@@
The formula (3.8) is used to compute the exact values of sensitivities $\rho_{j}$.
We proceed with estimation of (3.7). Since from (2.4) and (2.5), we have
$\|\mathbf{H}_{M}(0)\|_{2}$ $\leq\|\mathbf{V}_{M}(\boldsymbol{z})\|_{2}\,\|\mathbf{\Gamma}_{M}\|_{2}\,\|\mathbf{V}^{T}_{M}(\boldsymbol{z})\|_{2},$
@@
and taking into account the fact $\|\mathbf{X}\,\boldsymbol{e}_{j}\|_{2}\leq\|\mathbf{X}\|_{2}$, we obtain
@@
Thus
@@
Note that we can proceed further with the estimation of (3.10) applying known results for upper bounds of condition numbers of the Vandermonde matrices [15, 2].
Define further a vector of unstructured sensitivities $\boldsymbol{\rho}:=(\rho_{1},\ldots,\rho_{M})^{T}\in\mathbb{R}^{M}$ and let $\boldsymbol{\zeta}:=(\zeta_{1},\ldots,\zeta_{M})^{T}\in\mathbb{R}^{M}$. Then from (3.10) for the $\ell_{2}$-norm of the vector $\boldsymbol{\rho}$ we get
@@
From (3.8), taking into account (3.9) and (3.11), we obtain
@@
By summing up the last inequalities over $j=1,\ldots,M$, we get
$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}$
where $\zeta_{\max}:=\max_{j=1,\ldots,M}\zeta_{j}$. Finally using the property $\sum_{j=1}^{M}\|\mathbf{X}\boldsymbol{e}_{j}\|_{2}^{2}=\|\mathbf{X}\|_{\mathrm{F}}^{2}$, we obtain

### v3

@@ Abstract
In this paper, we introduce a new approach for the recovery of rational functions. The concept we propose is based on using the exponential structure of the Fourier coefficients of rational functions and the reconstruction of this exponential structure in the frequency domain. We choose ESPRIT as a method for the exponential recovery. The matrix pencil structure of this approach is the reason for its selection, as it makes our method suitable for the sensitivity analysis. According to our method, poles located inside and outside the unit circle are reconstructed independently as eigenvalues of some special Hankel matrix pencils. Furthermore, we derived formulas for sensitivities of poles of rational functions in case of unstructured and structured perturbations. Finally, we consider several numerical experiments and, using sensitivities, explain the recovery errors for poles.
@@ 1 Introduction
For $M\in\mathbb{N}$, we define a rational function $r(z)$ of type (M1,M)(M-1,M) by means of its partial fraction decomposition
@@
where $\gamma_{j},\,z_{j}\in\mathbb{C}$, coefficients $\gamma_{j}\neq 0$ and poles $z_{j}$ are pairwise distinct. We study the question of recovery of rational functions (1.1) using as input data a finite set of their Fourier coefficients $\widehat{r}_{k}$ defined by (4.1). We use the property that Fourier coefficients of rational functions have the exponential structure, and we reconstruct this exponential structure in the frequency domain applying the ESPRIT method [20, 18]. The idea of recovery of rational functions by estimation of signal parameters in the frequency domain is not new. Similar technique for the recovery of or approximation by rational functions was employed in [7, 21, 22, 23]. The paper [22] is of special interest to us. In [22], the author considered recovery of rational functions (1.1) based on using noisy input data and discussed reconstruction results for the poles $z_{j}$, j=1,,Mj=1,...,M. It was pointed out that for different levels of noise, not all poles $z_{j}$ are reconstructed with the same accuracy, and sometimes even the locations of certain poles are not reconstructed correctly. From these numerical observations, we can conclude that some poles $z_{j}$ of $r$ as in (1.1) are more ”sensitive“ to perturbations in the input data than the others. Our goal in this paper is to present theoretical explanations of these numerical observations by applying the sensitivity analysis. The method developed in [22] employees the AAK theory [17] for the reconstruction of the exponential structure of the Fourier coefficients $\widehat{r}_{k}$. We modify these ideas, by applying the ESPRIT method in the frequency domain. Since ESPRIT is based on using the Hankel pencil, it makes our method suitable for the sensitivity analysis.
@@
In [24], the authors developed the sensitivity analysis for the eigenvalues of Loewner pencil with respect to the unstructured and structured perturbations. We refer the interested reader to [1] for further application of Loewner framework for model reduction of dynamical systems. Following the approach in [24], we present the sensitivity analysis for the poles of $r$ in (1.1) computed by our method. According to our idea, poles located inside and outside the unit circle are reconstructed separately as eigenvalues of some Hankel matrix pencils. Further, we consider two kinds of Hankel pencil perturbations. For the sensitivity $\rho$ with respect to unstructured perturbations, we present the connection with the condition numbers of Vandermonde matrices that are used in factorizations of the corresponding Hankel matrices. The sensitivity $\eta$ with respect to structured perturbations of Hankel pencil is studied in [25]. We present modification of these results for our special rational reconstruction problem (1.1). Note that the theory of sensitivities of eigenvalues of Hankel pencil perturbations was considered also in [12, 3]. In [12] and [3], the authors measured both structured and unstructured perturbations by the same quantity. Similarly as in [24], we introduce two different quantities, sensitivities $\eta$ and $\rho$ with respect to structured und unstructured perturbations, respectively, for poles of the rational function (1.1). The ill-disposedness of $z_{j}$ is measured by determining whether $\rho$ and $\eta$ are ”large“ for this specific pole $z_{j}$. Additionally, the sensitivity $\eta$ reveals the perturbations of which Fourier coefficient $\widehat{r}_{k}$ the pole $z_{j}$ is most sensitive to.
@@
The method developed in this paper is also closely related to our ESPIRA method for exponential recovery [9][11] (see also [8] for the multivariate case). ESPIRA uses the property that the classical Fourier coefficients of exponential sums have a rational structure and reconstructs this rational structure with the AAA method for rational approximation [15]. Thus, the rational reconstruction concept employed in this paper can be viewed as the opposite of the ESPIRA concept. We would also like to draw particular attention to the paper [5], in which the authors used the sensitivity analysis of Hankel pencil eigenvalues to developed a method for exponential reconstruction.
@@
Finally, we want to mention the other methods for the recovery of rational functions such as the AAA method [15], the RKFIT (Rational Krylov Fitting) method [4], the Vector Fitting algorithm [14] and the Loewner framework method [1]. In [13], the authors discussed all these algorithms for the recovery of rational functions using as input information given noisy set of samples of $r$. Since we are interested in the sensitivity analysis of the poles, we have to assume that we apply perturbations to the Fourier coefficients $\widehat{r}_{k}$. Note that in this paper, we also introduce the corresponding modification of our method, which is based on employing function values of $r$ by computing $\widehat{r}_{k}$ via a quadrature formula for some grid of points (see Remark 4.1). Modifying our algorithm in this way, allows us to compare its performance with known rational methods. Numerical experiments show us that our approach gives equally good and, for some examples, even better recovery results from given noisy function values than the AAA method.
@@
Notation. As usual, $\mathbb{N}$, $\mathbb{Z}$, $\mathbb{R}$ and $\mathbb{C}$ denote the natural, integer, real and complex numbers, respectively. Let us agree that throughout the paper, we use the matrix notation $\mathbf{X}_{L,K}:=(X_{i,j})_{i,j=1}^{L,K}$ for matrices of the size $L\times K$ and the submatrix notation $\mathbf{X}_{L,K}(m:n,k:\ell)$, to denote a submatrix of $\mathbf{X}_{L,K}$ with rows indexed from $m$ to $n$ and columns indexed from $k$ to $\ell$, where first row and first column has index 1 (even though the row and column indices for the definition of the matrix may start with 0). For square matrices, we often use the short notation $\mathbf{X}_{K}$ instead of $\mathbf{X}_{K,K}$. By $\mathbf{X}_{L,K}^{\dagger}$ we denote the Moore-Penrose inverse of $\mathbf{X}_{L,K}$ and by $\mathbf{X}_{K}^{-1}$ inverse of the invertible matrix $\mathbf{X}_{K}$. We consider $\|\mathbf{X}_{L,K}\|_{2}:=\sigma_{\max}(\mathbf{X}_{L,K})$ the 2-norm or spectral norm, and $\|\mathbf{X}_{L,K}\|_{\mathrm{F}}:=\left(\sum_{i=1}^{L}\sum_{j=1}^{K}|X_{i,j}|^{2}\right)^{1/2}=\left(\sum_{i=1}^{\min\{L,K\}}\sigma_{i}^{2}\right)^{1/2}$ the Frobenius norm of the matrix $\mathbf{X}_{L,K}$. By $\boldsymbol{x}=(x_{1},\ldots,x_{K})^{T}$ we denote a column vector of the size $K$. The $\ell_{1}$- and $\ell_{2}$-norms of a vector $\boldsymbol{x}$ are given by $\|\boldsymbol{x}\|_{1}=\sum_{j=1}^{K}|x_{j}|$ and $\|\boldsymbol{x}\|_{2}=\left(\sum_{j=1}^{K}|x_{j}|^{2}\right)^{1/2}$, respectively.
@@ 2 ESPRIT method for the recovery of exponential sums
As previously mentioned in Introduction, we employ the ESPRIT method [20, 18] to reconstruct the exponential structure of the Fourier coefficients of rational functions (1.1). In this section, we describe the key concepts of this approach.
Let $f$ be an exponential sum of order $M\in\mathbb{N}$ defined by
$f(t):=\sum\limits_{j=1}^{M}\gamma_{j}z_{j}^{t},\ \gamma_{j},z_{j}\in\mathbb{C},$ (2.1)
where nodes $z_{j}$ are pairwise distinct and coefficients $\gamma_{j}\neq 0$. Let $L,N\in\mathbb{N}$ be given such that $M\leq L\leq N$, and $L$ is the upper bound for $M$. Our goal is to reconstruct the exponential sum (2.1), i.e. to reconstruct parameters $M$, $\gamma_{j}$ and $z_{j}$ for $j=1,\ldots,M$, using as input data a vector of samples $\boldsymbol{f}:=\big(f(0),f(1),\ldots,f(2N-1)\big)^{T}$ of the size $2N$.
@@
Since Vandermonde matrices in (2.2) have full column ranks, the matrix $\mathbf{H}_{2N-L,L+1}$ has rank $M$. Thus, in case of exact given data, the order $M$ of the exponential sum (2.1) is reconstructed as the numerical rank of the matrix $\mathbf{H}_{2N-L,L+1}$.
@@
From the last formula, we conclude that nodes $z_{j}$, $j=1,\ldots,M$ in (2.1) are eigenvalues of the matrix pencil (2.3). In case when $M$ is known, we can choose $N=L=M$ and compute $z_{j}$ as eigenvalues of
@@
Next, we discuss how we solve the matrix pencil (2.3). Consider SVD of ${\mathbf{H}}_{2N-L,L+1}$,
@@
where ${\mathbf{U}}_{2N-L}$ and ${\mathbf{W}}_{L+1}$ are unitary matrices and, in case of unperturbed data,
@@
where $\sigma_{1}\geq\ldots\geq\sigma_{M}>0$ are singular values of ${\mathbf{H}}_{2N-L,L+1}$. Consequently, (2.7) can be simplified to
@@
where the last L+1ML+1-M zero-columns in ${\mathbf{D}}_{2N-L,L+1}$ and the last L+1ML+1-M zero-rows in ${\mathbf{W}}_{L+1}$ are removed. Similarly as before, SVDs of ${\mathbf{H}}_{2N-L,L}(0)$ and ${\mathbf{H}}_{2N-L,L}(1)$ can be directly derived from (2.8) and the matrix pencil (2.3) now reads
@@
Therefore, eigenvalues of the matrix pencil (2.3) can be computed as eigenvalues of the matrix pencil (2.9). Then we define ${\mathbf{W}}_{M,L}(0)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,1\!:\!L)$ and ${\mathbf{W}}_{M,L}(1)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,2\!:\!L+1)$, and determine the vector $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ as a vector of eigenvalues of $\left({\mathbf{W}}_{M,L}(0)^{T}\right)^{\dagger}{\mathbf{W}}_{M,L}(1)^{T}$.
Having nodes $z_{j}$, we compute coefficients $\gamma_{j}$, $j=1,\ldots,M$ as a least squares solution of the system of linear equations
@@ 3 Sensitivities of Hankel pencil eigenvalues
The concept of sensitivities arises from the perturbations of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [12] and further considered in [3]. In [24], the notion of sensitivities was defined for Loewner matrix pencil. In the paper [24], the authors define two sensitivities, which correspond to structured and unstructured perturbations. Using similar ideas, in [25], the author considered the structured sensitivity of Hankel matrix pencil. In [5], the sensitivity analysis of Hankel pencil was applied to develop a new method for the exponential analysis. In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3) and in Subsection 4.2, we present modification of these results to develop sensitivity analysis for the poles of rational functions (1.1).
3.1 Sensitivities of Hankel pencil eigenvalues with unstructured perturbations
Following the approach for Loewner pencil from [24], we show that in case of unstructured perturbations of Hankel pencil, the sensitivities of eigenvalues depend on the numerical conditions of Vandermonde matrices from factorizations (2.4) and (2.5) (for $N=L=M$). We consider the generalized eigenvalue problem (GEP) for the Hankel matrix pencil (2.6)
@@ 3.1 Sensitivities of Hankel pencil eigenvalues with unstructured perturbations
where $\boldsymbol{p}=(p_{1},\ldots,p_{M})^{T}$ and $\boldsymbol{q}=(q_{1},\ldots,q_{M})^{T}$ are left and right eigenvectors, respectively, of the GEP (3.1). Then under the perturbations
$\tilde{\mathbf{H}}_{M}(0):=\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)},\ \ \tilde{\mathbf{H}}_{M}(1):=\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)},$ (3.2)
@@
Formula (3.3) is derived from the perturbed system
@@
Assume that the $\ell_{2}$-norms of perturbation matrices $\Delta_{\mathbf{H}_{M}(0)}$ and $\Delta_{\mathbf{H}_{M}(1)}$ in (3.2) are bounded as follows
@@
for some $\varepsilon>0$. Then from (3.3) taking into account (3.4), we obtain
@@
where $\rho$ is the sensitivity with respect to unstructured perturbations of the eigenvalue $z$ of the GEP (3.1) defined by
@@
Next, we derive formulas for the exact computation of sensitivities $\rho_{j}$ of the eigenvalues $z_{j}$, $j=1,\ldots,M$ of the GEP (3.1) as well as their upper bounds. Since the Hankel matrices $\mathbf{H}_{M}(0)$ and $\mathbf{H}_{M}(1)$ satisfy factorizations (2.4) and (2.5) for $N=L=M$, left $\boldsymbol{p}^{(j)}$ and right $\boldsymbol{q}^{(j)}$ eigenvectors corresponding to the eigenvalue $z_{j}$ can be determined by formulas
@@
where $\boldsymbol{e}_{j}$ is the $j$th canonical column vector in $\mathbb{R}^{M}$. Then using again the factorization (2.4) of the Hankel matrix $\mathbf{H}_{M}(0)$, we get
@@
From (3.5) taking into account (3.7), we obtain that the sensitivity $\rho_{j}$ is given by
@@
The ill-disposedness of the eigenvalue $z_{j}$ with respect to unstructured perturbations can be measured by checking whether $\rho_{j}$ is ”large“. Applying (2.4), (2.5) and (3.6) to (3.8), we obtain the following final formula for the computation of sensitivities $\rho_{j}$
@@
We proceed with the estimation of (3.8). Since (2.4) and (2.5) implies
$\|\mathbf{H}_{M}(0)\|_{2}$ $\leq\|\mathbf{V}_{M}(\boldsymbol{z})\|_{2}\,\|\mathbf{\Gamma}_{M}\|_{2}\,\|\mathbf{V}^{T}_{M}(\boldsymbol{z})\|_{2},$ (3.10)
@@
and taking into account the inequality $\|\mathbf{X}\,\boldsymbol{e}_{j}\|_{2}\leq\|\mathbf{X}\|_{2}$, we conclude that
@@
Thus, we derive the following upper bounds for sensitivities $\rho_{j}$,
@@
Note that we can estimate (3.12) further by applying known estimates for condition numbers of Vandermonde matrices [16, 2].
Defining $\boldsymbol{\zeta}:=(\zeta_{1},\ldots,\zeta_{M})^{T}\in\mathbb{R}^{M}$ and the vector of sensitivities $\boldsymbol{\rho}:=(\rho_{1},\ldots,\rho_{M})^{T}\in\mathbb{R}^{M}$, and applying (3.12), we get for the $\ell_{2}$-norm of the vector $\boldsymbol{\rho}$,
@@
From (3.9), taking into account (3.10), (3.11) and (3.13), we obtain
@@
By summing up the last inequalities over $j=1,\ldots,M$, we arrive at
$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big\|\mathbf{V}_{M}(\boldsymbol{z})\big\|_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\bigg\|_{2}^{2},$
where $\zeta_{\max}:=\max_{j=1,\ldots,M}\zeta_{j}$. Finally, using the property $\sum_{j=1}^{M}\|\mathbf{X}\boldsymbol{e}_{j}\|_{2}^{2}=\|\mathbf{X}\|_{\mathrm{F}}^{2}$, we derive the following upper bound